OpenStudy (anonymous):

Solve the quadratic equation by completing the square: x^2 -8x + 13 = 0

OpenStudy (anonymous):

-8x*.5 = -4, (-4)^2 = 16, x^2 -8x + 16 = -13 (x-4)^2 = -13 Then I square root both sides but my calculator says I can't square root -13

Parth (parthkohli):

That means that this has no real solution...

Parth (parthkohli):

Wait, there is..

Parth (parthkohli):

Yes, right

Parth (parthkohli):

Come to twid with me, and I'll check

OpenStudy (anonymous):

Oops, x = 4 +- sqrt(3)

OpenStudy (anonymous):

Twid?